OpenStudy (anonymous):

Evaluate log7 136 Round your answer to the nearest thousands

OpenStudy (anonymous):

So we're starting with the following: \[\log_7 136=x\]The equation we'll be using is this: \[\log_n x = \ln x / \ln n\]So our problem becomes: \[\frac{ \ln 136 }{ \ln 7 }\]Which you can leave as is, or simplify as needed.

OpenStudy (anonymous):

2.52460 is the answer. Rounding it off to the nearest thousandths will give you 2.525.